Solve the triangle.


A = 52°, b = 14, c = 6


1
Expert's answer
2021-06-14T19:02:31-0400

Answer:


B=103.8° , a=11.3 and C=24.2°


Given:

A = 52°, b = 14, c = 6


using sine law:

"{a\over sinA}={b\over sinB}={c\over sinC}"


"{a\over sin52^o}={14\over sinB}={6\over sinC}" ..........................1


using cosine law:


"a^2=b^2+c^2-2bcCosA"


"a^2=196+36-2(14)(6)Cos52^o" .............2

from 2 we get,

a=11.3 ........................3


putting 3 in 1

And solving

"{11.3\over sin52^o}={14\over sinB}"

And

"{11.3\over sin52^o}={6\over sinC}"

As we see first possible solution

we get

B=103.80

C=24.20